OpenStudy (anonymous):

9x + 5y =-20 and -4x -3y=12

OpenStudy (ageta):

x=0 and y=−4

OpenStudy (ageta):

I will try to solve your system of equations. 9x+5y=−20;−4x−3y=12 Step: Solve 9x+5y=−20 for x: 9x+5y+−5y=−20+−5y(Add -5y to both sides) 9x=−5y−20 /9 /9 x=−5/9 y + -20/9

OpenStudy (ageta):

\[\left(\begin{matrix}-5 \\ 9\end{matrix}y+ -20/9\right) for -x =-4x-3y=12:-4x-3y=12\]